METHANE LINE PARAMETERS IN THE NEAR-INFRARED

\underline{L. R. BROWN}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

To characterize the methane spectrum from 4800-5500 and 6180 to 10000 cm^{−1}, over 40000 positions and line intensities have been retrieved from selected laboratory spectra. These data have been recorded at room temperature and at 0.01 and 0.02 cm^{−1} resolution using the McMath-Pierce Fourier transform spectrometer located at Kitt Peak, AZ. Positions have been calibrated using residual water features at 5300 and 7200 cm^{−1}. Intensity measurements are good to 7% or better for the stronger features; weaker absorptions are less accurately measured. These new measurements have been listed in HITRAN format and mixed with the old information for the 5500 to 6180 cm^{−1} interval. Probable assignments from the literature have been listed for 2% of the observed features.

a

aThe research described in this paper was carried out by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.